negative 16 and one third divided by (negative 2 and one third)
Answer by Alan3354(69443)

negative 16 and one third divided by (negative 2 and one third)
= (-49/3)/(-7/3)
the -1's cancel
= (49/3)/(7/3)
Multiply NUM and DEN by 3
= (49)/(7)
= 7
---------
Or -
(-49/3)/(-7/3)
Invert the 2nd and multiply
= (49/3)*(3/7)
= 147/21
= 7